OpenStudy (anonymous):

help? Solve 8^x - 2 = 3 x ≈ 2.52 x ≈ 0.26 x ≈ 2.74 x ≈ 0.40

OpenStudy (compassionate):

Hello. \[8^x - 2 = 3 \rightarrow 8^x = 5\] Now just take the root of 8 to solve.

OpenStudy (anonymous):

c?

OpenStudy (blurbendy):

are you sure those are the correct answer choices?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

x should be something like .77 though...

OpenStudy (anonymous):

wait, I think you got the question wrong..kind of. its 8^(x - 2 )= 3

OpenStudy (blurbendy):

ooo..

OpenStudy (anonymous):

so?

OpenStudy (blurbendy):

8(2.52 - 2) ≈ 3

OpenStudy (anonymous):

so how do I do it..?

OpenStudy (blurbendy):

take the log base 8 of both sides log(base8)8 ^ (x - 2) = x-2 log(base8) 3 = log(3) / log(8) x - 2 = log(3) / log(8) x = 2 + log(3) / log(8)

OpenStudy (anonymous):

okay..i think im following..

OpenStudy (anonymous):

@Compassionate can you help???????

OpenStudy (compassionate):

The opposite of 8 to the x is the logarithm of 8. Which is?

OpenStudy (anonymous):

I have no idea..

OpenStudy (compassionate):

The logaritm of 8 is 2.

OpenStudy (anonymous):

oh..

OpenStudy (compassionate):

So take it to both sides log(based: 8) = 5 -(log base 8)

OpenStudy (anonymous):

okay

OpenStudy (compassionate):

So then we get x = 0.74

OpenStudy (anonymous):

yup got it. thanks!